OpenStudy (anonymous):

Determine whether the graphs of the given pair of lines are perpendicular y=4x-5 4y=8-x Please show the work, so I can figure out how to do this. If someone can explain it simply, it would be appreciated. Thank you.

OpenStudy (anonymous):

slopes are 4 and -1/4 product is -1 lines are perpendicular

OpenStudy (anonymous):

Thank you, but how did you reach those answers?

OpenStudy (anonymous):

y = 4x -5 compare with y=mx+b m=4 4y=-x+ 8 y=-x/4 +2 m = -1/4

OpenStudy (anonymous):

Thank you.. I am catching on to it now.. I appreciate the help :)

OpenStudy (anonymous):

welcome:)